# Experimental confirmation of the magnetic ordering transition induced by an electronic structure change in the metallic triangular antiferromagnet Co1/3TaS2

### v1

@@ Abstract
We report ARPES studies combined with DFT+DMFT calculations to confirm that the magnetic ordering vector transition from Q=(1/2,0,0) to Q=(1/3,0,0) in the metallic triangular antiferromagnets Co1/3±ϵTaS2 (0.007) is induced by the electronic structure change in the system. The ARPES-measured Fermi surface (FS) maps of Co0.325TaS2 show two hexagonal and one circular hole-like FSs around , which matches well with the triple-Q state by taking into account the contribution of nesting vectors occurring between Co 3 and Ta 5 orbitals. In the case of Co0.340TaS2, a new electron pocket around K appears and the FS geometry changes as a result of the correlation effect of Co4S18 tripods forming in the system. The magnetic susceptibility calculations based on the DFT+DMFT band structure indicate that the most stable magnetic ordering vector changes to (1/3,0,0) from (1/2,0,0), which is very consistent with the magnetic phase transition around =1/3 in CoxTaS2.
This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-public-access-plan).
@@
In this paper, employing synchrotron-based ARPES combined with density functional theory plus dynamic mean field theory (DFT+DMFT), we closely scrutinized the changes in the electronic structure of Co1/3±ϵTaS2 (0.007) single crystals having a minute difference in cobalt content to understand the origin of the magnetic ordering vector transition and the details for stabilization of the 3Q magnetic ordering. The ARPES measurements reveal that the Fermi surface geometry changes drastically in spite of the small deviation from 1/3 moles of cobalt ions per formula unit. It is definitely beyond electron filling effect in the rigid band picture, and can be explained only as a result of correlation effect induced by the slightly over-doped cobalt ions in the Co0.340TaS2 samples. The magnetic susceptibility calculations based on the DFT+DMFT band structures suggest that the FS geometry change is closely connected with the magnetic ordering vector transition from Q=(1/2,0,0) to Q=(1/3,0,0) in reciprocal lattice units even though our DMFT spin susceptibility calculation does not include the momentum-dependent effects of self-energy and vertex function. They also imply that the scatterings between Co 3 and Ta 5 electrons are the key process for the 3Q ordering stabilization. These results demonstrate that although it is evident in theory but hard to verify experimentally, a local spin ordering transition can be induced solely through changes in the electronic structure of a metallic magnet.
@@
CoxTaS2 is Co-intercalated 2-TaS2, where the magnetic Co ions form a ordered phase in the vdW gap when =1/3.[18] The unit cell of the crystal structure is shown in Fig. 1(a), in which a Co ion is located at the center of a S6 octahedron with D3h symmetry. The basic transport properties were measured in the 1980s to show a metallic behavior with an antiferromagnetic transition at 35 K, and a strong magnetic phase dependent Hall coefficient.[19] However, the recent revisited measurements on the single crystals revealed that the system with 0.30 0.325 has another magnetic phase below 26.5 K to show a significant anomalous Hall effect (AHE), which is strongly coupled to the weak ferromagnetism.[20, 21, 22] The exact magnetic structure for this phase was recently solved only after combining powder/single crystal neutron diffraction and AHE measurements, giving the tetrahedral 3Q ordering with Q=(1/2,0,0) in reciprocal lattice units.[11] The 3Q state with the AHE shows a sensitive dependence on the Co content as shown in the magnetic phase diagram of Fig. 1(b).[21] This sensitivity provides a rare opportunity to investigate what stabilizes the 3Q state and how it does so. Although our previous ARPES measurements give a hint to the electronic origin for the magnetic transition, a close examination of combined experiment and theory for the electronic structure on both sides of the phase boundary can give the quantitative understanding for the stabilization mechanism. Two different Co composition samples, =0.325 (hereafter, UD-CTS) and 0.340 (OD-CTS), were chosen to measure the electronic structure by ARPES. The Co content difference is just 0.015 moles per formula unit, so effectively there is no difference in Co ions arrangement except the existence of Co4S18 tripods in OD-CTS as shown in Fig. 1(c). The density of the tripods in this work is smaller than 0.7 % per formula unit.
@@
The drastic change of the electronic structure from UD-CTS to OD-CTS induced by the small amount of additional Co intercalants is definitely beyond a simple Fermi level tuning effect, and we argue that it is qualitatively explained by introducing electron correlation effects mediated by Co4S18 tripods in OD-CTS. An additional Co intercalant more than 1/3 moles per formula unit occupies a vacant Co site with =c/4 or 3c/4 (the 2 Wyckoff position in the space group ) in a Co layer as illustrated in Fig. 1(c), which forms a Co4S18 tripod.[20] Once the system hosts Co4S18 tripods, they are expected to induce additional interaction channels in the electronic structure and in the magnetic structure, respectively. Since a Co atom occupies the center of a trigonal anti-prismatic S6 octahedron with symmetry, the Co 3 orbitals are split into three manifolds e, a1g, and e as shown in Fig. 4(a). A simple perturbative estimate of the crystal-field energy splitting, based on the atomic positions obtained from diffraction refinements,[20] yields an energy difference between a1g and e states of approximately 0.31 of the total crystal field splitting . Then, a divalent Co cat-ion with seven 3 electrons has half-filled e and a1g states with a high spin configuration =3/2. The distance between the nearest Co ions in a tripod is 3.31 Å, which is the second shortest bonding in the system.[21] At this distance, direct overlap between Co 3 orbitals is non-negligible. Within a tripod, the central Co 3 orbitals hybridize equally with those of the three corner Co atoms, with a coupling strength . Because hybridization among the corner Co sites is negligible, the Co cluster in the tripod forms one bonding, two non-bonding, and one anti-bonding state per 3 orbital, resulting in an energy gain of . If the bonding energy lowering exceeds the crystal-field splitting , an electron from the e level transfers into the bonding a1g state, as illustrated in Fig. 4(b). Another electron from the e states may occupy the non-bonding a1g level depending on the energy of the anti-bonding e states. In the band structure, this direct bonding leads to partially filled bands with predominat Co 3 a1g character, which explains the appearance of the band electron pocket at the K point. This argument is also consistent with the observation that the change in the electronic structure appears dominantly only in the bands. Since these are derived from Co - interactions, bands with strong Co 3 characters are most strongly affected. As is shown in Fig. 3(a1), all conduction bands exhibit enhanced Co contributions near M and K points, while the Co 3 weight is reduced near point. Because the conduction band minimum of the band is located at K, it is qualitatively parallel with the observed behavior.
@@
In order to theoretically confirm the dominancy of the electronic structure effect to the spin ordering pattern, we performed the band structure calculations based on density functional theory plus dynamical mean field theory (DFT+DMFT), and computed the magnetic susceptibility based on the DMFT band structures. The detailed calculation methods are described in Calculational Methods. Figures 5(a)-(c) display the DMFT band structure, the DMFT FS, and the magnetic susceptibility calculations for UD-CTS, respectively. To simulate the band structure of UD-CTS, we use the Hubbard interaction =2 eV, the Hund’s coupling =0.4 eV in the calculation. Even though the value is somewhat smaller than the typically used value in transition metal ions, it gives an optimal result for our ARPES-measured band structure of CoxTaS2 within a reasonable parameter range.[16, 17] The Fermi level was adjusted by introducing an additional doping parameter =0.8, which means an additional number of valence electrons per crystallographic unit cell, i.e. per two Co cations. This additional doping in DMFT is necessary to compare with the surface-sensitive ARPES measurements. The DMFT band structure shows the renormalization of the DFT one (green line) near the Fermi energy. If we compare the conduction bands in the DMFT FS with the experimental ones, the calculated band is smaller than the experimental one. However, the major features of UD-CTS band structure are qualitatively well reproduced in the calculation. The momentum dependency of the magnetic susceptibility was also checked as shown in Fig. 5(c) using the eq. (1) and (2) described in Calculational Methods. While the polarizability has the weak momentum dependence, the full susceptibility shows the strong momentum dependence, which is enhanced by the increase of the RPA-type inter-orbital interaction between the Co 3 and Ta 5 orbitals. For UD-CTS, the leading peak position is at M, indicating that the nesting of Q=(1/2,0,0) occurs between the Co 3 ( band) and Ta 5 ( band) orbital characters as denoted with red arrow in Fig. 5(b). For OD-CTS, we increase the additional doping parameter to 1.2, which moves parts of the band structure near K towards below the Fermi level. This is clearly seen by comparing the calculated bands denoted with a white dotted ellipse in Fig. 5(a) and (d), respectively. The difference =0.4 in the calculations corresponds to additional electron doping of 0.2 per Co ion, while the electron doping difference between OD-CTS and UD-CTS in the experiments is 0.042 electrons per Co ion. This large difference clearly indicates that the FS changes are not induced by a simple Fermi level shift in rigid bands, and electron correlation effects from the tripods have to be taken into account. As is described above, the Co-Co direct bondings in the tripods play a key role to push down the energy states with relatively strong Co 3 orbital character near the zone boundary including K point. Unfortunately, the effect of the low density tripods and the short-range correlation between Co ions can not be fully incorporated in the DFT+DMFT scheme, so the deviation between the calculations and the ARPES data is inevitable to some degree. For example, the spectral intensity near K in Fig. 5(e) is much weaker than that of ARPES data.
@@
In spite of the inevitable deviation, the calculated magnetic susceptibility well explains the magnetic ordering vector of Q=(1/3,0,0) in the over-doped region as shown in Fig. 5(f). In the magnetic susceptibility calculations, it is worth mentioning that not only the electron doping but also the interplay between the FS nesting vector and the RPA-type interactions plays an important role to the transition of the magnetic ordering vector. Although the change of is not too sensitive to the doping, the nesting effect based on multiple peaks can be clearly seen as in Fig. 5(g). While the leading peak position is located at the M point due to the inter-band nesting effect of the Ta 5 hole pockets (red arrow in Fig. 5(b)), the second highest peak, which is possibly responsible for the helical order in the over-doped region, is located near Q=(1/3,0,0) due to the nesting effect within Co 3 hole pockets (additional red arrow in Fig. 5(e)). Upon the electron doping, the peak near Q=(1/3,0,0) is more strongly enhanced than the other comparable peak at Q=(1/2,0,0), and it becomes the leading peak at =1.2 due to the large RPA-type intra-atomic Co 3 interaction as shown in Fig. 5(f). Even though the exact leading peak position is a little smaller than 1/3 of a reciprocal lattice constant (the vertical dotted line between -M), this matches well with the nesting vector within the Co 3 bands, indicating that the FS geometry dominates the magnetic ordering patterns in Co1/3+ϵTaS2.
@@
In conclusion, we performed ARPES studies combined with DFT+DMFT calculations to confirm that the magnetic ordering vector transition in Co1/3±ϵTaS2 (0.007) is induced by an electronic structure change. The ARPES-measured FS map of UD-CTS has two hexagonal and one circular hole-like Fermi surfaces around , which matches well with the 3Q state by taking into account the contribution of nesting vectors occurring between Co 3 and Ta 5 orbitals. In OD-CTS case, a new electron pocket around K appears and the FS geometry changes as a result of the correlation effect of Co4S18 tripods. The magnetic susceptibility calculations based on the DFT+DMFT band structure indicate that the most stable magnetic ordering vector changes to approximately Q=(1/3,0,0) from Q=(1/2,0,0), which is very consistent with the magnetic phase transition around =1/3 in CoxTaS2.
I Experimental Details
II Calculational Methods
We adopt density functional theory plus dynamical mean field theory (DFT+DMFT) to study the correlated band structure and the magnetic susceptibility of Co1/3TaS2. To perform DFT+DMFT, we first obtain the non-spin-polarized band structure using the Vienna Ab-initio Simulation Package (VASP) code. We used the experimental crystal structure of Co1/3TaS2, i.e. the ordered phase by Co intercalants. The Perdew-Burke-Ernzerhof (PBE) functional is used for the exchange-correlation energy of DFT. In DFT, we also used a 14144 -mesh with the energy cutoff of 400 eV for the plane-wave basis. Using the DFT non-spin-polarized band structure, we construct the tight-binding Hamiltonian that can reproduce effectively the same band structure to the DFT calculations, using the maximally localized Wannier functions for Co 3 and Ta 5 orbitals. Since other Ta 5 orbitals are located above the Fermi level, including them in the tight binding Hamiltonian is not essential. This Hamiltonian is solved using DMFT by adopting the continuous-time quantum Monte Carlo (CTQMC) method as the impurity solver. We use the Hubbard interaction =2 eV, the Hund’s coupling =0.4 eV for Co 3 orbitals, and the temperature T=116 K. While the adopted value is smaller than the reported value obtained by the contrained-LDA method, it is within a reasonable range.[17] In DMFT, we used a fine -mesh of 303010. For the double counting potential, we used the fully-localized-limit formula of the double counting potential. The number of valence electrons is set to be 24.0. To impose the doping effect, we change the number of valence electrons and rerun the DMFT calculation for the fixed number of valence electrons. To compute the magnetic susceptibility, we use the following formula based on the DMFT Green’s function, as explained in Ref.[16]. First, we compute the polarizability as follows:
III Acknowledgments

### v2

@@ Abstract
We report ARPES studies combined with DFT+DMFT calculations to confirm that the magnetic ordering vector transition from Q=(1/2,0,0) to Q=(1/3,0,0) in the metallic triangular antiferromagnets Co1/3±ϵTaS2 (0.007) is induced by the electronic structure change in the system. The ARPES-measured Fermi surface (FS) maps of Co0.325TaS2 show two hexagonal and one circular hole-like FSs around , which matches well with the triple-Q state by taking into account the contribution of nesting vectors occurring between Co 3 and Ta 5 orbitals. In the case of Co0.340TaS2, a new electron pocket around K appears and the FS geometry changes as a result of the correlation effect of Co4S18 tripods forming in the system. The magnetic susceptibility calculations based on the charge-self-consistent DFT+DMFT band structures and the random phase approximation indicate that the most stable magnetic ordering vector (1/2,0,0) split into (1/6,0,0) and (1/2,0,0), which is consistent with the magnetic phase transition around =1/3 in CoxTaS2.
I Introduction
@@
In this paper, employing synchrotron-based ARPES combined with density functional theory plus dynamic mean field theory (DFT+DMFT), we closely scrutinized the changes in the electronic structure of Co1/3±ϵTaS2 (0.007) single crystals having a minute difference in cobalt content to understand the origin of the magnetic ordering vector transition and the details for stabilization of the 3Q magnetic ordering. The ARPES measurements reveal that the Fermi surface geometry changes drastically in spite of the small deviation from 1/3 moles of cobalt ions per formula unit. It is definitely beyond electron filling effect in the rigid band picture, and can be explained only as a result of correlation effect induced by the slightly over-doped cobalt ions in the Co0.340TaS2 samples. The magnetic susceptibility calculations based on the DFT+DMFT band structures and the random phase approximation (RPA) suggest that the FS geometry change is closely connected with the magnetic ordering vector transition from Q=(1/2,0,0) to Q=(1/3,0,0) in reciprocal lattice units even though our DMFT spin susceptibility calculation does not include the momentum-dependent effects of self-energy and vertex function. They also imply that the scatterings between Co 3 and Ta 5 electrons are the key process for the 3Q ordering stabilization. These results demonstrate that although it is evident in theory but hard to verify experimentally, a local spin ordering transition can be induced solely through changes in the electronic structure of a metallic magnet.
II Experimental Details
III Calculational Methods
We adopt density functional theory plus dynamical mean field theory (DFT+DMFT) to study the correlated band structure and the magnetic susceptibility of Co1/3TaS2. To perform DFT+DMFT, we first obtain the non-spin-polarized band structure using the Vienna Ab-initio Simulation Package (VASP) code. We used the experimental crystal structure of Co1/3TaS2, i.e. the ordered phase by Co intercalants. The Perdew-Burke-Ernzerhof (PBE) functional is used for the exchange-correlation energy of DFT. In DFT, we also used a 14144 -mesh with the energy cutoff of 400 eV for the plane-wave basis. Using the DFT non-spin-polarized band structure, we construct the tight-binding Hamiltonian that can reproduce effectively the same band structure to the DFT calculations, using the maximally localized Wannier functions for Co 3 and Ta 5 orbitals. Since other Ta 5 orbitals are located above the Fermi level, including them in the tight binding Hamiltonian is not essential. This Hamiltonian is solved using DMFT by adopting the continuous-time quantum Monte Carlo (CTQMC) method as the impurity solver. We use the Hubbard interaction =3 eV, the Hund’s coupling =0.4 eV for Co 3 orbitals, and the temperature T=116 K. While the adopted value is smaller than the reported value obtained by the contrained-LDA method, it is within a reasonable range.[19] We achieved the charge-self-consistency of DFT+DMFT using the DMFTwDFT code.[20] In DMFT, we used a fine -mesh of 303010. For the double counting potential, we used the fully-localized-limit formula of the double counting potential. The number of valence electrons is set to be 24.0. To impose the doping effect, we change the number of valence electrons and rerun the DMFT calculation for the fixed number of valence electrons. To compute the magnetic susceptibility, we use the following formula based on the DMFT Green’s function, as explained in Ref.[16]. First, we compute the polarizability as follows:
IV Results
@@
CoxTaS2 is Co-intercalated 2-TaS2, where the magnetic Co ions form a ordered phase in the vdW gap when =1/3.[21] The unit cell of the crystal structure is shown in Fig. 1(a), in which a Co ion is located at the center of a S6 octahedron with D3h symmetry. The basic transport properties were measured in the 1980s to show a metallic behavior with an antiferromagnetic transition at 35 K, and a strong magnetic phase dependent Hall coefficient.[22, 23] However, the recent revisited measurements on the single crystals revealed that the system with 0.30 0.325 has another magnetic phase below 26.5 K to show a significant anomalous Hall effect (AHE), which is strongly coupled to the weak ferromagnetism.[24, 17, 25] The exact magnetic structure for this phase was recently solved only after combining powder/single crystal neutron diffraction and AHE measurements, giving the tetrahedral 3Q ordering with Q=(1/2,0,0) in reciprocal lattice units.[11] The 3Q state with the AHE shows a sensitive dependence on the Co content as shown in the magnetic phase diagram of Fig. 1(b). For the detailed spin configurations for the each magnetic phase, see Figs. 4(f)-(h) in Ref. [17]. This sensitivity provides a rare opportunity to investigate what stabilizes the 3Q state and how it does so. Although our previous ARPES measurements give a hint to the electronic origin for the magnetic transition, a close examination of combined experiment and theory for the electronic structure on both sides of the phase boundary can give the quantitative understanding for the stabilization mechanism. Two different Co composition samples, =0.325 (hereafter, UD-CTS) and 0.340 (OD-CTS), were chosen to measure the electronic structure by ARPES. The Co content difference is just 0.015 moles per formula unit, so effectively there is no difference in Co ions arrangement except the existence of Co4S18 tripods in OD-CTS as shown in Fig. 1(c). The density of the tripods in this work is smaller than 0.7 % per formula unit.
@@
The drastic change of the electronic structure from UD-CTS to OD-CTS induced by the small amount of additional Co intercalants is definitely beyond a simple Fermi level tuning effect, and we argue that it is qualitatively explained by introducing electron correlation effects mediated by Co4S18 tripods in OD-CTS. An additional Co intercalant more than 1/3 moles per formula unit occupies a vacant Co site with =c/4 or 3c/4 (the 2 Wyckoff position in the space group ) in a Co layer as illustrated in Fig. 1(c), which forms a Co4S18 tripod.[24] Once the system hosts Co4S18 tripods, they are expected to induce additional interaction channels in the electronic structure and in the magnetic structure, respectively. Since a Co atom occupies the center of a trigonal anti-prismatic S6 octahedron with symmetry, the Co 3 orbitals are split into three manifolds e, a1g, and e as shown in Fig. 4(a). A simple perturbative estimate of the crystal-field energy splitting, based on the atomic positions obtained from diffraction refinements,[24] yields an energy difference between a1g and e states of approximately 0.31 of the total crystal field splitting . Then, a divalent Co cat-ion with seven 3 electrons has half-filled e and a1g states with a high spin configuration =3/2. The distance between the nearest Co ions in a tripod is 3.31 Å, which is the second shortest bonding in the system.[17] At this distance, direct overlap between Co 3 orbitals is non-negligible. Within a tripod, the central Co 3 orbitals hybridize equally with those of the three corner Co atoms, with a coupling strength . Because hybridization among the corner Co sites is negligible, the Co cluster in the tripod forms one bonding, two non-bonding, and one anti-bonding state per 3 orbital, resulting in an energy gain of . If the bonding energy gain exceeds the on-site Coulomb energy cost in e or a1g states, electrons in the e or levels form spin-singlet bonding states, as illustrated in Fig. 4(b). In the band structure, this direct bonding leads to a down-shift of Co character bands, which explains the appearance of the band electron pocket at the K point. This argument is also consistent with the observation that the change in the electronic structure appears dominantly only in the bands. Since these are derived from Co - interactions, bands with strong Co 3 characters are effectively affected. As is shown in Fig. 3(a1), all conduction bands exhibit enhanced Co contributions near M and K points, while the Co 3 weight is reduced near point. Because the conduction band minimum of the band is located at K, it is qualitatively parallel with the observed behavior.
@@
In order to theoretically confirm the dominancy of the electronic structure effect to the spin ordering pattern, we performed the band structure calculations based on charge-self-consistent (CSC) density functional theory plus dynamical mean field theory (DFT+DMFT), and computed the magnetic susceptibility based on the DMFT band structures. The detailed calculation methods are described in Calculational Methods. Figures 5(a)-(c) display the DMFT band structure, the DMFT FS, and the magnetic susceptibility calculations for UD-CTS, respectively. To simulate the band structure of UD-CTS, we use the Hubbard interaction =3 eV, the Hund’s coupling =0.4 eV in the calculation. Even though the value is somewhat smaller than the typically used value in transition metal ions, it gives an optimal result for our ARPES-measured band structure of CoxTaS2 within a reasonable parameter range.[16, 19] The Fermi level was adjusted by introducing an additional doping parameter =0.8, which means an additional number of valence electrons per crystallographic unit cell, i.e. per two Co cations. This additional doping in DMFT is necessary to compare with the surface-sensitive ARPES measurements. The DMFT band structure shows the renormalization of the DFT one (green line) near the Fermi energy. If we compare the conduction bands in the DMFT FS with the experimental ones, the calculated band is smaller than the experimental one. However, the major features of UD-CTS band structure are qualitatively well reproduced in the calculation. The momentum dependency of the magnetic susceptibility was also checked as shown in Fig. 5(c) using the eq. (1) and (2) described in Calculational Methods. While the polarizability has the weak momentum dependence, the full susceptibility shows the strong momentum dependence, which is enhanced by the increase of the RPA-type inter-orbital interaction between the Co 3 and Ta 5 orbitals. For UD-CTS, the leading peak position is at M, indicating that the nesting of Q=(1/2,0,0) occurs between the Co 3 ( band) and Ta 5 ( band) orbital characters as denoted with red arrow in Fig. 5(b). For OD-CTS, we increase the additional doping parameter to 1.2, which moves parts of the band structure near K towards below the Fermi level. This is clearly seen by comparing the calculated bands denoted with a white dotted ellipse in Fig. 5(a) and (d), respectively. The difference =0.4 in the calculations corresponds to additional electron doping of 0.2 per Co ion, while the electron doping difference between OD-CTS and UD-CTS in the experiments is 0.042 electrons per Co ion. This large difference clearly indicates that the FS changes are not induced by a simple Fermi level shift in rigid bands, and electron correlation effects from the tripods have to be taken into account. As is described above, the Co-Co direct bondings in the tripods play a key role to push down the energy states with relatively strong Co 3 orbital character near the zone boundary including K point. Unfortunately, the effect of the low density tripods and the short-range correlation between Co ions can not be fully incorporated in the DFT+DMFT scheme, so the deviation between the calculations and the ARPES data is inevitable to some degree. For example, the spectral intensity near K in Fig. 5(e) is much weaker than that of ARPES data.
@@
In spite of the inevitable deviation, the calculated magnetic susceptibility explains the change of the magnetic ordering vector Q=(1/2,0,0) in the over-doped region as shown in Fig. 5(f). In the magnetic susceptibility calculations, it is worth mentioning that not only the electron doping but also the interplay between the FS nesting vector and the RPA-type interactions plays an important role to the transition of the magnetic ordering vector. Although the change of is not too sensitive to the doping, the nesting effect based on multiple peaks can be clearly seen as in Fig. 5(g). While the leading peak position is located at the M point due to the inter-band nesting effect of the Ta 5 hole pockets (red arrow in Fig. 5(b)), the second highest peak, which is possibly responsible for the helical order in the over-doped region, is located near Q=(1/3,0,0) due to the nesting effect within Co 3 hole pockets (additional red arrow in Fig. 5(e)). Upon the electron doping, the peak near Q1/3=(1/3,0,0) moves to Q1/6=(1/6,0,0) and is more strongly enhanced than the other comparable peak at Q1/2=(1/2,0,0) to become the leading peak at =1.2 due to the large RPA-type intra-atomic Co 3 interaction as shown in Fig. 5(f). This definitely makes Q1/2 unstable and Q1/6 more stable for the magnetic ordering vector. However, the peak weight at Q1/2 is still very comparable to the new leading peak at Q1/6, so the resulting magnetic ordering vector is not simply predictable. The experimentally determined ordering vector Q1/3 is probably a competing result between the two Q vectors. Even though the exact leading peak position in OD-CTS is not clearly derived from the magnetic susceptibility calculations, this matches well with the nesting vector within the Co 3 bands, indicating that the FS geometry dominates the magnetic ordering patterns in Co1/3+ϵTaS2.
V Conclusion
@@
In conclusion, we performed ARPES studies combined with CSC-DFT+DMFT calculations to confirm that the magnetic ordering vector transition in Co1/3±ϵTaS2 (0.007) is induced by an electronic structure change. The ARPES-measured FS map of UD-CTS has two hexagonal and one circular hole-like Fermi surfaces around , which matches well with the 3Q state by taking into account the contribution of nesting vectors occurring between Co 3 and Ta 5 orbitals. In OD-CTS case, a new electron pocket around K appears and the FS geometry changes as a result of the correlation effect of Co4S18 tripods. The magnetic susceptibility calculations based on the DFT+DMFT band structures and the RPA indicate that the most stable magnetic ordering vector Q=(1/2,0,0) split into Q=(1/6,0,0) and (1/2,0,0), which is consistent with the magnetic phase transition around =1/3 in CoxTaS2.
VI Acknowledgments